# X-ray panorama of the SS 433/W50 complex by SRG/eROSITA

### v1

@@ 3 Broad-band imaging
We now turn to the deeper PV data. The broad-band (0.5-4 keV) vignetting-corrected X-ray image of the entire W50 nebula is shown in Fig. 2, with the contours of radio emission at 4.85 GHz (Gregory et al., 1996) overlaid in white. The color coding reflects the measured X-ray surface brightness on a square-root scale spanning two orders of magnitude, highlighting the wide dynamic range of the diffuse X-ray emission captured by eROSITA. In addition to SS 433 itself (the extremely bright source in the center) and a multitude of point sources, the majority of which are nearby foreground stars or background AGN, the picture shows bright diffuse emission with rich internal structure.
The X-ray picture can be broadly decomposed into three components. Firstly, there is a low surface brightness emission filling the entire projected extent of the nebula. It is brighter in the Eastern and Western lobes of W50 (left and right, respectively, in Fig. 2) and significantly dimmer along the Northern and Southern radio boundaries of W50. Remarkably, the central part of W50 (beyond the region contaminated by the PSF wings from SS 433) appears even fainter in X-rays, suggesting a shell-like geometry of the X-ray-emitting gas.
@@
Secondly, one can notice the presence of very narrow X-ray filaments, particularly inside and close to the boundaries of the Eastern and Western lobes. Their apparent width is , while their length is . Given the angular resolution of eROSITA (better than , i.e., at 5 kpc), these filaments are fully resolved in both directions. The X-ray filaments form a complicated network with an overall structure very similar to the filaments of the cold and dense gas observed via H+[N II] emission at different locations within the nebula (Boumis et al., 2007; Farnes et al., 2017). The largest X-ray filaments have a flat transverse profile of the surface brightness and are an order of magnitude brighter than the surrounding medium. Such filaments of the dense X-ray emitting gas might arise as instabilities in the sheared flows, caustics of the strong shock waves, or due to heating of the preexisting cold gas filaments by hot gas and relativistic particles inside the lobes.
@@
The final and perhaps the most visual feature of the image is a pair of nearly symmetrical, sharp-edged “extended X-ray jets” directed away from the central source. Their axis is not only aligned with the elongation of the radio nebula and location of the X-ray bright lobes, but also coincides with the projection of the precession axis of SS 433’s famous trans-relativistic baryonic jets. The physical nature of these structures, first found by the Einstein observatory (Watson et al., 1983) and then observed with all major X-ray telescopes (Brinkmann et al., 1996; Safi-Harb & Ögelman, 1997; Namiki et al., 2000; Brinkmann et al., 2007; Safi-Harb & Petre, 1999), including the most recent polarimteric observations with IXPE (Kaaret et al., 2024), remains a matter of debate (e.g. Churazov et al., 2024). The SRG/eROSITA data clearly show a very rapid onset of emission at the innermost boundary of EXJs on both sides of the nebula (also demonstrated recently by Tsuji et al., 2025, using Chandra data).
4 Spectrally-decomposed imaging
@@
We take advantage of the unique possibility provided by /eROSITA to probe the spectral characteristics of diffuse X-ray emission throughout the full extent of the nebula for the first time. To produce a spectrally-decomposed image, we divided the full spectral band presented earlier into three sub-bands: soft (0.5-1 keV), medium (1-2 keV), and hard (2-4 keV) bands, respectively. A composite RGB image showing the surface brightness of the emission (now on a linear scale) in these bands in red, green, and blue, respectively, is presented in Figure 3. The chosen spectral decomposition immediately reveals the presence of two distinct components in the observed diffuse emission: large regions of red-orange color where emission below 1 keV dominates, and green-cyan structures where prominent emission above 1 keV is present (including the bright central source SS 433). Compact sources also form two groups: the softer (“red-yellow”) ones are predominantly foreground stars, while harder (“blueish”) sources are mostly background AGNs, which shine through the gas and dust in the disk of the Milky Way. Zoom-in images of the lobes (rotated and put side by side to help the comparison) are shown in Fig. 4 and demonstrate fine structure of the soft diffuse emission, contrasting with relatively smooth and well-confined hard emission.
@@
For better characterization of the diffuse emission, we identified bright point sources and masked them along with the central source SS 433. The resulting composite map of the residual diffuse emission is shown in Figure 5. The overlaid distance and angle rulers highlight the characteristic dimensions and morphology of the prominent structures. Namely, one can see that the “Extended X-ray Jets” start pc away from the central source and demonstrate remarkable symmetry, with the Eastern one starting only a couple of parsecs further from the central source. Both EXJs, characterized by hard X-ray spectra, terminate around pc. This symmetry contrasts with the more asymmetric and irregular soft emission of the lobes.
@@
The same conclusion is even better illustrated by the side-by-side comparison of the two lobes in Fig. 4. Indeed, while the Eastern lobe extends up to , the Western one terminates at pc. The inclination of the relativistic jet’s precession axis to the line of sight is (the relativistic jet on the eastern side is on average approaching us), so this difference cannot be explained by geometrical effects only. However, it broadly agrees with the assumption that the energy flow from the central source interacts with the denser ambient medium in the western part, reflecting the gas density gradient perpendicular to the equatorial plane of the Galactic disk (e.g., Goodall et al., 2011a). Since the difference between the lobes’ dimensions amounts to a sizeable fraction of them, one can infer the exponential scale-height of the ambient gas density profile pc, consistent with the scale-height of the Galactic neutral gas disk. The X-ray emission from the Northern and Southern boundaries of the nebula appears soft and faint, being noticeably affected by the foreground dust absorption, especially in their Western parts (as illustrated in Appendix C).
@@
Taking this into account, we conclude that the soft X-ray emission can be traced across the full extent of the nebula, except for the “inner cavity” with a radius of pc. The outer boundary of the emission of the Northern and Southern arcs (outlined as dashed annuli including Region 4 in Fig. 5) is located at from the center, coinciding with the boundary of the W50’s radio emission. This morphology suggests that the soft X-ray emitting gas is confined in a thick () close-to-spherical shell, driving a strong shock in the surrounding ISM.
@@
The Eastern and the Western lobes are roughly confined inside the projection of the cone with the half-opening angle , while the EXJs encompass only the cone with a twice smaller opening angle,. From the eROSITA spectrally decomposed RGB images, a gentle spectral evolution of the EXJs emission with distance is seen clearly. EXJ emission is significantly harder at its sharp inner boundary than further away from SS 433, where EXJs blend with softer thermal emission that fills the rest of the lobe, in agreement with earlier findings with XMM-Newton and NuSTAR (Brinkmann et al., 2007; Safi-Harb et al., 2022).
@@ 5.1 Central source
The spectral capabilities of eROSITA have been demonstrated using many objects, including SS 433 itself (e.g., Sunyaev et al., 2021). Since this source is extremely bright, the observations were intentionally performed during the two consecutive orbital eclipses of the compact object (and the jets) by the companion star, so its soft X-ray flux was dimmer by a factor of several compared to the typical levels (Kawai et al., 1989; Lopez et al., 2006; Marshall et al., 2013). The combined X-ray spectrum of SS 433 extracted from arcmin aperture (with background estimated from to 5 arcmin annuli, being however completely negligible compared the source) using the data of all pointings in which it falls within Field-of-View is shown in Fig. 6.
As expected, it is well described by the model of the optically thin multitemperature emission from collimated flows that adiabatically expand and cool at distances from a few to a few cm from the central object (Brinkmann et al., 1988, 1991, 2005; Kotani et al., 1996; Marshall et al., 2002; Medvedev & Fabrika, 2010; Khabibullin et al., 2016; Medvedev et al., 2019). We discuss it more in Appendix A, while the main focus of this study is instead on the much fainter diffuse emission from the surrounding nebula.
@@ 5.2 Diffuse emission
The high sensitivity of the presented data allowed us to perform spectroscopy of the diffuse emission in several regions and determine the physical state of the radiating plasma. We have selected eight regions, shown as numbered boxes in Figure 5, for detailed spectral analysis of their X-ray emission. For each region, the background (estimated from the adjacent regions) was subtracted to isolate the net signal. We present the parameters of the fitted models in Table 1 and discuss the implications of the corresponding results below.
@@ 5.2.1 Eastern Lobe
Emission from the brightest filament in the Eastern lobe (Region 1 in Figure 5) features a spectrum characteristic for a thermal (but non-equilibrium) optically-thin plasma with a temperature of , as evidenced by the line emission from the highly ionized oxygen, neon, magnesium, silicon and iron (see the left panel in Figure 7). The line emissivity and ratios are consistent with the Solar abundance of heavy elements, but also indicate a non-equilibrium state of the gas ionization. Namely, it is qualitatively consistent with the expectation for a plane-parallel shell of the gas that experienced a passage of the shock with velocity (from the Rankine-Hugoniot relations for the temperature and density jump properties) some yr ago.
The inferred hydrogen number density, cm-3, is a factor of a few higher than the expected gas density of the interstellar gas pc away from the plane of the Galactic disk. The spectrum of the diffuse emission in the Eastern lobe (exemplified by Region 3 in Figure 5) also can be well described by thermal non-equilibrium emission of the gas with somewhat higher temperature keV and lower hydrogen number density cm-3 (see the left panel in Figure 7).
@@ 5.2.2 Western Lobe
The emission from the Western lobe appears to be noticeably harder, partially due to the higher interstellar absorption in that direction. However, it also corresponds to the emission of plasma with genuinely higher temperature, keV, and a similar hydrogen number density, cm-3 (and a few times higher for the brightest filament), as illustrated by spectra from Regions 6 and 7 in Figure 5). The non-equilibrium ionization model (see the right panel in Figure 7) results in a comparable estimate of yr for the ionization time-scale.
The total absorption-corrected luminosity of the diffuse soft X-ray emission from the nebula is in 0.5-4 keV, which is by a factor of few tens smaller than the apparent X-ray luminosity of the central object, few ,222Due to the beaming of the radiation along the thick accretion disc axis (expected in the case of highly supercritical accretion), the actual X-ray luminosity of SS 433 might be orders of magnitude larger, namely comparable to the Eddington luminosity and kinetic luminosity of the jets, (e.g. Fabrika & Mescheryakov, 2001; Begelman et al., 2006; Poutanen et al., 2007; Medvedev & Fabrika, 2010; Khabibullin & Sazonov, 2016; Middleton et al., 2021; Fogantini et al., 2023). and much smaller than the supposed kinetic luminosity of the outflows from the supercritical accretion disk, . However, the thermal energy content of the X-ray emitting gas in the Eastern lobe (with estimated gas number density 0.1 cm-3, temperature 0.5 keV and volume assuming roughly to cylindrical geometry ) is erg, requiring a supply of at least over years. The mass of this gas amounts to , implying that the bulk of it cannot be provided by the outflows from SS 433. If the volume filling factor of the emitting medium is, however, low, the estimate for the required mass and energy could be decreased by a factor of , but similarly, the density should be increased, leading to the corresponding decrease in the time since the shock passage. That means that the X-ray emitting gas provides an estimate of the average energy input in the lobes at the level of , which is comparable to the kinetic luminosity of the jets and the Eddington luminosity of the system (for the mass of the compact object below 10).
A similar estimate can be obtained from the Western lobe, implying nearly symmetric energy injection along the nebula’s axis at a very high rate over the last few tens of thousands of years. Given that the temperature inferred from the spectral fits with the non-equilbirum emission is driven mostly by characteristics of the ionization state, it is likely a lower limit on the actual temperature of the emitting gas. Hence, a significant portion of the gas energy content might be effectively hidden, while the required energy input should be increased correspondingly.
@@ 5.2.3 Northern and southern boundaries
The emission from the nebula boundaries is exemplified by the Northern Arc region and the Southern Extension (Regions 4 and 8 in Figure 5). The thermal non-equilibrium model provides the temperature estimate keV, low hydrogen number density, cm-3 for the Northern Arc and even lower for the Southern Extension. The ionization timescale is, however, similar here as well, yr. This observation contrasts with the brightness of the radio emission, which in the Northern Arc region is one of the brightest across the nebula. The high degree of linear polarization of the radio emission might indicate that the preferential orientation of the interstellar magnetic field in this region favors efficient acceleration of relativistic particles producing synchrotron emission. We note here that the central part of the W50 X-ray map suffers from higher photoelectric absorption than the Eastern and even Western lobes, as illustrated in Fig. 9. This excess absorption is likely associated with the foreground gas and has no physical connection to SS 433/W50.

### v2

@@ 3 Broad-band imaging
We now turn to the deeper PV data. The broad-band (0.5-4 keV) vignetting-corrected X-ray image of the entire W50 nebula is shown in Fig. 2, with the contours of radio emission at 4.85 GHz (Gregory et al., 1996) overlaid in white. The color-coding reflects the measured X-ray surface brightness on a square-root scale spanning two orders of magnitude, highlighting the wide dynamic range of the diffuse X-ray emission captured by eROSITA. In addition to SS 433 itself (the extremely bright source in the center) and a multitude of point sources –the majority of which are nearby foreground stars or background active galactic nuclei (AGNs) – the picture shows bright diffuse emission with rich internal structure.
@@
The X-ray picture can be broadly decomposed into three components. First, there is a low-surface-brightness emission filling the entire projected extent of the nebula. It is brighter in the eastern and western lobes of W50 (left and right, respectively, in Fig. 2) and significantly dimmer along the northern and southern radio boundaries of W50. Remarkably, the central part of W50 (beyond the region contaminated by the PSF wings from SS 433) appears even fainter in X-rays, suggesting a shell-like geometry of the X-ray-emitting gas. Second, one can notice the presence of very narrow X-ray filaments, particularly inside and close to the boundaries of the eastern and western lobes. Their apparent width is , while their length is . Given the angular resolution of eROSITA (better than ; i.e., at 5 kpc), these filaments are fully resolved in both directions. The X-ray filaments form a complicated network with an overall structure very similar to the filaments of the cold and dense gas observed via H+[N II] emission at different locations within the nebula (Boumis et al., 2007; Farnes et al., 2017). The largest X-ray filaments have a flat transverse profile of the surface brightness and are an order of magnitude brighter than the surrounding medium. Such filaments of the dense X-ray-emitting gas might arise as instabilities in the sheared flows, caustics of the strong shock waves, or due to heating of the preexisting cold-gas filaments by hot gas and relativistic particles inside the lobes.
@@
The final and perhaps the most visual feature of the image is a pair of nearly symmetrical, sharp-edged EXJs directed away from the central source. Their axis is not only aligned with the elongation of the radio nebula and location of the X-ray bright lobes, it also coincides with the projection of the precession axis of SS 433’s famous trans-relativistic baryonic jets. The physical nature of these structures, first found by the Einstein observatory (Watson et al., 1983) and then observed with all major X-ray telescopes (Brinkmann et al., 1996; Safi-Harb and Ögelman, 1997; Namiki et al., 2000; Brinkmann et al., 2007; Safi-Harb and Petre, 1999), including the most recent polarimteric observations with IXPE (Kaaret et al., 2024), remains a matter of debate (e.g., Churazov et al., 2024). The /eROSITA data clearly show a very rapid onset of emission at the innermost boundary of EXJs on both sides of the nebula (also demonstrated recently using Chandra data by Tsuji et al., 2025).
4 Spectrally decomposed imaging
@@
We took advantage of the unique possibility provided by /eROSITA to probe the spectral characteristics of diffuse X-ray emission throughout the full extent of the nebula for the first time. To produce a spectrally decomposed image, we divided the full spectral band presented earlier into three subbands: soft (0.5-1 keV), medium (1-2 keV), and hard (2-4 keV) bands, respectively. A composite pseudo-color (red-green-blue) image showing the surface brightness of the emission (now on a linear scale) in these bands in red, green, and blue, respectively, is presented in Figure 3. The chosen spectral decomposition immediately reveals the presence of two distinct components in the observed diffuse emission: large regions in red-orange where emission below 1 keV dominates, and green-cyan structures where prominent emission above 1 keV is present (including the bright central source SS 433). Compact sources also form two groups: the softer (red-yellow) ones are predominantly foreground stars, while harder (blueish) sources are mostly background AGNs, which shine through the gas and dust in the disk of the Milky Way. Zoomed-in images of the lobes (rotated and put side by side to aid comparison) are shown in Fig. 4 and demonstrate fine structure of the soft diffuse emission, contrasting with relatively smooth and well-confined hard emission.
@@
For better characterization of the diffuse emission, we identified bright point sources and masked them along with the central source SS 433. The resulting composite map of the residual diffuse emission is shown in Figure 5. The overlaid distance and angle rulers highlight the characteristic dimensions and morphology of the prominent structures. Namely, one can see that the EXJs’ start pc away from the central source and demonstrate remarkable symmetry, with the eastern one starting only a couple of parsecs further from the central source. Both EXJs, characterized by hard X-ray spectra, terminate at around pc. This symmetry contrasts with the more asymmetric and irregular soft emission of the lobes.
@@
The same conclusion is even better illustrated by the side-by-side comparison of the two lobes in Fig. 4. Indeed, while the eastern lobe extends up to , the western one terminates at pc. The inclination of the relativistic jet’s precession axis to the line of sight is (the relativistic jet on the eastern side is on average approaching us), so this difference cannot be explained by geometrical effects alone. However, it broadly agrees with the assumption that the energy flow from the central source interacts with the denser ambient medium in the western part, reflecting the gas density gradient perpendicular to the equatorial plane of the Galactic disk (e.g., Goodall et al., 2011a). Since the difference between the lobes’ dimensions amounts to a sizeable fraction of them, one can infer the exponential scale height of the ambient gas density profile of pc, which is consistent with the scale height of the Galactic neutral-gas disk. The X-ray emission from the northern and southern boundaries of the nebula appears soft and faint and is noticeably affected by the foreground dust absorption, especially in their western parts (as illustrated in Appendix C).
@@
Taking this into account, we conclude that the soft X-ray emission can be traced across the full extent of the nebula, except for the “inner cavity,” with a radius of pc. The outer boundary of the emission of the northern and southern arcs (outlined as dashed annuli including Region 4 in Fig. 5) is located at from the center, coinciding with the boundary of W50’s radio emission. This morphology suggests that the soft X-ray-emitting gas is confined in a thick () close-to-spherical shell, driving a strong shock in the surrounding ISM.
@@
The eastern and western lobes are roughly confined inside the projection of the cone with the half-opening angle of , while the EXJs only encompass the cone with an opening angle that is two times smaller, i.e., (see the characteristic angles indicated in Fig. 5). The opening angle of the baryonic jets launched by the central source is much smaller: (e.g., Marshall et al., 2002). Therefore, if the jet direction precesses slowly and the ejected matter moves ballistically, it would only fill thin “walls” of a hollow cone, making a tightly wound spiral with cm pc separation ( at kpc). Due to a light-travel-time delay between the approaching and receding parts of this spiral, however, the apparent spacing between individual twists should be completely smeared out at distances beyond 1 pc of the central source (e.g., Khabibullin and Sazonov, 2017). From the eROSITA spectrally decomposed pseudo-color (red-green-blue) images, a gentle spectral evolution of the EXJs’ emission with distance is seen clearly. EXJ emission is significantly harder at its sharp inner boundary than further away from SS 433, where EXJs blend with softer thermal emission that fills the rest of the lobe; this is in agreement with earlier findings of XMM-Newton and NuSTAR (Brinkmann et al., 2007; Safi-Harb et al., 2022).
@@ 5.1 Central source
The spectral capabilities of eROSITA have been demonstrated using many objects, including SS 433 itself (e.g., Sunyaev et al., 2021). Since this source is extremely bright, the observations were intentionally performed during the two consecutive orbital eclipses of the compact object (and the jets) by the companion star, so its soft X-ray flux was dimmer by a factor of several compared to the typical levels (Kawai et al., 1989; Lopez et al., 2006; Marshall et al., 2013). The combined X-ray spectrum of SS 433 extracted from an arcmin aperture (with background estimated from to 5 arcmin annuli; this is, however, completely negligible compared the source) using the data of all pointings in which it falls within the field of view is shown in Fig. 6.
As expected, it is well described by the model of the optically thin multi-temperature emission from collimated flows that adiabatically expand and cool at distances from a few to a few cm from the central object (Brinkmann et al., 1988, 1991, 2005; Kotani et al., 1996; Marshall et al., 2002; Medvedev and Fabrika, 2010; Khabibullin et al., 2016; Medvedev et al., 2019). We discuss this further in Appendix A; the main focus of this study is instead on the much fainter diffuse emission from the surrounding nebula.
@@ 5.2 Diffuse emission
The high sensitivity of the presented data allowed us to perform spectroscopy of the diffuse emission in several regions and determine the physical state of the radiating plasma. We selected eight regions, shown as numbered boxes in Figure 5, for a detailed spectral analysis of their X-ray emission. For each region, the background (estimated from the adjacent regions) was subtracted to isolate the net signal. We present the parameters of the fit models in Table 4 and discuss the implications of the corresponding results below.
@@ 5.2.1 Eastern lobe
Emission from the brightest filament in the eastern lobe (Region 1 in Figure 5) features a spectrum characteristic for a thermal (but nonequilibrium) optically-thin plasma with a temperature of , as evidenced by the line emission from the highly ionized oxygen, neon, magnesium, silicon, and iron (see the left panel in Figure 7). The line emissivity and ratios are consistent with the solar abundance of heavy elements, but they also indicate a nonequilibrium state of the gas ionization. In particular, the simultaneous presence of spectral lines from ion species correspond to a broad range of gas temperatures (from helium-like oxygen O VII to numerous iron L-shell ions), likely reflecting an ongoing ionization process of initially cold and rapidly heated ISM gas (although a multi-temperature mixture of the equilibrium spectra is also capable of reproducing similar features). The obtained best-fit, nonequillibrium spectral model is qualitatively consistent with the expectation for a plane-parallel shell of the gas that experienced a passage of the shock with velocity, (from the Rankine-Hugoniot relations for the temperature and density jump properties) yr ago (taking mean molecular weight ).
The inferred hydrogen number density, cm-3, is a factor of a few higher than the expected gas density of the interstellar gas pc away from the plane of the Galactic disk. The spectrum of the diffuse emission in the eastern lobe (exemplified by Region 3 in Figure 5) also can be well described by thermal nonequilibrium emission of the gas with a somewhat higher temperature of keV and lower hydrogen number density of cm-3 (see the left panel in Figure 7).
@@ 5.2.2 Western lobe
The emission from the western lobe appears to be noticeably harder, partially due to the higher interstellar absorption in that direction. However, it also corresponds to the emission of plasma with genuinely higher temperature, keV, and a similar hydrogen number density, cm-3 (this is a few times higher for the brightest filament), as illustrated by spectra from Regions 6 and 7 in Figure 5). The nonequilibrium ionization model (see the right panel in Figure 7) results in a comparable estimate of yr for the ionization timescale.
The total absorption-corrected luminosity of the diffuse soft X-ray emission from the nebula is in 0.5-4 keV, which is several tens times smaller than the apparent X-ray luminosity of the central object, few ,222Due to the beaming of the radiation along the thick-accretion-disk axis (expected in the case of highly supercritical accretion), the actual X-ray luminosity of SS 433 might be orders of magnitude larger and namely comparable to the Eddington luminosity and kinetic luminosity of the jets: (e.g., Fabrika and Mescheryakov, 2001; Begelman et al., 2006; Poutanen et al., 2007; Medvedev and Fabrika, 2010; Khabibullin and Sazonov, 2016; Middleton et al., 2021; Fogantini et al., 2023). and much smaller than the supposed kinetic luminosity of the outflows from the supercritical accretion disk: . However, the thermal-energy content of the X-ray-emitting gas in the eastern lobe (with an estimated gas number density of 0.1 cm-3, temperature of 0.5 keV, and volume of , assuming roughly cylindrical geometry) is erg, requiring a supply of at least over years. The mass of this gas amounts to , implying that the bulk of it cannot be provided by the outflows from SS 433. If the volume filling factor, of the emitting medium is low, however, the estimate for the required mass and energy could be decreased by a factor of , but similarly, the density should be increased, leading to the corresponding decrease in the time since the shock passage. That means that the X-ray-emitting gas provides an estimate of the average energy input in the lobes at the level of , which is comparable to the kinetic luminosity of the jets and the Eddington luminosity of the system (for the mass of the compact object below 10).
A similar estimate can be obtained from the western lobe, implying nearly symmetric energy injection along the nebula’s axis at a very high rate over the last few tens of thousands of years. Given that the temperature inferred from the spectral fits with the nonequilbirum emission is driven mostly by characteristics of the ionization state, it is likely a lower limit on the actual temperature of the emitting gas. Hence, a significant portion of the gas energy content might be effectively hidden, while the required energy input should be increased accordingly.
@@ 5.2.3 Northern and southern boundaries
The emission from the nebula boundaries is exemplified by the northern arc region and the southern extension (Regions 4 and 8 in Figure 5). The thermal-nonequilibrium model provides a temperature estimate of keV and low hydrogen number density of cm-3 for the northern arc, and this is even lower for the southern extension. The ionization timescale is, however, similar: yr. The derived parameters of the nonequillibrium ionization model for the northern-arc region (listed in Table 4) are in good agreement with the findings of Chi et al. (2024), which studied a similar region using the data of dedicated XMM-Newton observations guided by earlier mosaic of the whole W50 nebula by ROSAT PSPC (Brinkmann et al., 1996). A great advantage of the SRG/eROSITA data (particularly important for the very faint extended emission from these regions) is significantly more robust control over contributions of the instrumental and astrophysical backgrounds. This allows much tighter constraints to be obtained on the key parameters of interest here. The observation of rather faint X-ray emission contrasts with the brightness of the radio emission, which in the northern-arc region is one of the brightest across the nebula. The high degree of linear polarization of the radio emission might indicate that the preferential orientation of the interstellar magnetic field in this region favors efficient acceleration of relativistic particles producing synchrotron emission. We note that the central part of the W50 X-ray map suffers from higher photoelectric absorption than the eastern and even western lobes, as illustrated in Fig. 9. This excess absorption is likely associated with the foreground gas and has no physical connection to SS 433/W50.